A "moving sidewalk" in an airport terminal building moves at 1.0 m/s and is 35.0 m long. If a woman steps on at one end and walks at 1.5 m/s relative to the moving sidewalk, how much time does she require to reach the opposite and if she walks (a) in the same direction the sidewalk is moving? (b) In the opposite direction?

1
Expert's answer
2022-01-27T09:12:18-0500

a)

"t_1=\frac s{v+u}=14~s,"

b)

"t_2=\frac s{v-u}=70~s."